A business owner borrows $5,600 for 5 months. At the end of the 5-month loan period,
the interest owed is $115. What was the simple yearly interest rate
(as a percent)?

Answer by Boreal(15235)

I=PRT
R=I/PT
R=[115/(5600*5/12)]
R=115/(28000/12)=4.93%
Check with 5%, which should be close and a little higher
5600*(0.05)(5/12)=280*5/12=1400/12=116.67 which is consistent.